Solve the differential equation
xdy/dx +y = y^2Inx
1
Expert's answer
2020-12-04T07:31:07-0500

Given differential equation: "x\frac{dy}{dx}+y=y^2lnx"

Dividing the given equation by "x" , we get

"\frac{dy}{dx}+\frac{y}{x}=\frac{y^2lnx}{x}"

Observe that the above differential equation is Bernoulli's differential equation.

It can be reduced to Linear Differential equation dividing the equation by "y^2" and then take the substitution "\frac{1}{y}=z\Rightarrow \frac{1}{y^2}\frac{dy}{dx}=-\frac{dz}{dx}"

Then the given differential equation becomes

"\frac{1}{y^2}\frac{dy}{dx}+\frac{1}{y}\frac{1}{x}=\frac{lnx}{x}"

"\Rightarrow -\frac{dz}{dx}+\frac{z}{x}=\frac{lnx}{x}

\Rightarrow \frac{dz}{dx}-\frac{z}{x}=-\frac{lnx}{x}"

The above is a first order linear differential equation of the form

"\frac{dz}{dx}+p(x)z=g(x)" , where "p(x)=-\frac{1}{x}" and "g(x)=-\frac{lnx}{x}"

General solution to the differential equation is

"z" (Integrating Factor)"=\int" (Integrating Factor)("g(x)dx+C" , where

Integrating Factor "=e^{\int p(x)dx}"

Using the above, we have

Integrating Factor = "=e^{\int (-\frac{1}{x})dx}=e^{-lnx}=e^{ln(\frac{1}{x})}=\frac{1}{x}"

General solution is

"z(\frac{1}{x})=\int (\frac{1}{x})(-\frac{lnx}{x})dx+C=-\int \frac{lnx}{x^2}dx+C"

Take the substitution "lnx=u\Rightarrow \frac{1}{x}dx=du" and "\frac{1}{x}=e^{-u}"

Then, we get

"\frac{z}{x}=-\int (ue^{-u})du+C=-[-ue^{-u}-e^{-u}]+C" (Using integration by parts)

That is we get

"\frac{z}{x}=[ue^{-u}+e^{-u}]+C=\frac{lnx}{x}+\frac{1}{x}+C" "(u=lnx)"

That is we get

"\frac{1}{xy}=\frac{lnx}{x}+\frac{1}{x}+C(z=\frac{1}{y})"

Therefore, general solution to the given differential equation is

"ylnx+y+C=1"